# CogCoM: Train Large Vision-Language Models Diving into Details through Chain of Manipulations

### v2

@@ Abstract
Vision-Language Models (VLMs) have demonstrated their broad effectiveness thanks to extensive training in aligning visual instructions to responses. However, such training of conclusive alignment leads models to ignore essential visual reasoning, further resulting in failures in meticulous visual problems and unfaithful responses. Drawing inspiration from human cognition in solving visual problems (e.g., marking, zoom in), this paper introduces Chain of Manipulations, a mechanism that enables VLMs to solve problems step-by-step with evidence. After training, models can solve various visual problems by eliciting intrinsic manipulations (e.g., grounding, zoom in) with results (e.g., boxes, image) actively without involving external tools, while also allowing users to trace error causes. We study the roadmap to implement this mechanism, including (1) a flexible design of manipulations upon extensive analysis, (2) an efficient automated data generation pipeline, (3) a compatible VLM architecture capable of multi-turn multi-image, and (4) a model training process for versatile capabilities. With the design, we also manually annotate 6K high-quality samples for the challenging graphical mathematical problems. Our trained model, CogCoM, equipped with this mechanism with 17B parameters achieves state-of-the-art performance across 9 benchmarks from 4 categories, demonstrating the effectiveness while preserving the interpretability. Our code, model weights, and collected data will be publicly available.
Figure 1: CogCoM solves various visual problems with Chain of Manipulations mechanism. Note that the CoM reasoning generates evidential and explainable steps, without involving external tools.
@@ 1 Introduction
Benefiting from the advantage of Large Language Models (LLMs) in broad world knowledge, large Vision Language Models (VLMs) (Alayrac et al., 2022; Wang et al., 2023b) that are further trained to understand visual inputs have demonstrated viabilities on broad multimodal scenarios, such as visual question answering (Liu et al., 2023b), visual grounding (Peng et al., 2023), optical character recognition (Zhang et al., 2023b). The research employing VLMs as foundation models (Bai et al., 2023; Sun et al., 2023b; Wang et al., 2023b) usually involves two main stages of training, where the first stage develops intrinsic visual understanding ability through exposure to massive image-caption pairs, and the second stage endows the models with problem-solving capabilities through the instruction tuning.
@@
However, existing tuning methods train models to respond to instructions with conclusive language responses upon visual inputs, which leads models to ignore the essential intermediate visual reasoning and further results in failures in meticulous visual problems, unfaithful responses, and even hallucinations. For example in the left subplot of Figure 2, we test the top-performing model CogVLM (Wang et al., 2023b) about the details in the image (i.e., texts written on a pillar), and it directly responds an incorrect answer (i.e., NO SMOKING), most likely from bias to visual or linguistic priors (i.e., typical scenes with a pillar in office). The absence of the essential reasoning on the visual scene may lead to a rash response (Hwang et al., 2023).
@@
Humans solve problems regarding visual details by marking or processing the given images for convenience and rigor, which we refer to as manipulations. For example, we find targets by sequentially locating references, and concentrate on subtle details by zooming into a corresponding region. Most of VLMs have developed numerous intrinsic capabilities (e.g., grounding boxes, recognizing texts) during the first stage of training. By further imitating the fundamental human behaviours (e.g., cropping, zoom in), models have the potential to perform this cognitive reasoning process. Three major obstacles in eliciting VLMs with such reasoning are (1) flexible definitions of manipulations covering most visual problems, (2) an efficient data collection pipeline capable of producing abundant training data, and (3) a multi-turn multi-image VLM structure compatible with existing models.
@@
Inspired by the human cognition in solving visual problems, we introduce Chain of Manipulations (CoM), a mechanism that enables VLMs to solve problems step-by-step with evidence, with each step potentially involving a manipulation on the visual input and its corresponding result, both generated by the model to facilitate the success and fidelity. This paper studies a complete roadmap with manipulations design, data collection, model architecture and training process for training general VLMs with this mechanism. We first formally design 6 basic manipulations upon the pilot experiments, which are capable of handling diverse visual problems. Next, we propose a cascading data generation pipeline based on reliable large language models (e.g., LLMs, the linguistic annotators) and visual foundational models (e.g., VFMs, the visual annotators), which can automatically produce abundant error-free training data. We collect 70K CoM samples with this pipeline. We then devise a multi-turn multi-image model architecture compatible with typical VLMs structures. Based on a data recipe incorporating the curated corpus, we finally train a general VLM equipped with CoM reasoning mechanism, named CogCoM, which possesses capabilities of chat, captioning, grounding and reasoning. Additionally, benefiting from the expressive capability of the proposed mechanism, we further manually annotated 6K high-quality samples of graphical mathematical problems, each accompanied by a CoM reasoning process, to advance the research of VLMs in solving challenging mathematical problems.
@@
We conduct extensive experiments on 9 benchmarks from 4 categories, including TextVQA (Singh et al., 2019), ST-VQA (Biten et al., 2019), TallyVQA (Acharya et al., 2019), and GQA Hudson & Manning (2019) for detailed visual question answering, RefCOCO (Yu et al., 2016), RefCOCO+(Yu et al., 2016), and RefCOCOg (Mao et al., 2016) for visual grounding, POPE (Li et al., 2023c) for hallucination validation, and MM-Vet (Yu et al., 2023b) for general multimodal ability. Our model achieves up to 9.0 and 1.09 accuracy improvement on the detailed VQA and grounding benchmarks, respectively, and the superior performance on the general multimodal benchmark. The results demonstrate the effectiveness of the mechanism while maintaining the interpretability of outputs.
@@ 2 Terminology
We first conduct pilot experiments to investigate the possible manipulations capable of handling diverse visual problems.
@@
Specifically, given a question about an image, we prompt the advanced large language model, GPT-4, to generate solving steps by optionally utilizing possible actions on the image that facilitate problem-solving. We conduct this experiment on 170K questions from TextVQA, a dataset requiring detailed reasoning and recognition on images. To ensure the stability, we manually write 4 demonstrations as priors, The detailed statistics are available at Appendix C.3.
@@
Based on the observation, we formally predefine a set of 6 manipulations, which can either be developed from pre-training or be learned from fine-tuning with the imitation to human behaviors: $\mathcal{M}\subseteq$$\textit{OCR}(tgt)\rightarrow txt$, $\textit{Grounding}(tgt)\rightarrow bbx$, $\textit{Counting}(tgt)\rightarrow num$, $\textit{Calculate}(tgt)\rightarrow num$, $\textit{CropZoomIn}(bbx,x)\rightarrow img$, $\textit{Line}(pts)\rightarrow img$}, where the parameters or results $tgt,txt,bbx,num,x,img,pts$ refer to the bounding boxes, zoom ratio, image, target description, numbers, texts, and points, respectively. In addition to the predefined manipulations, we also allow trained models to create new manipulations during inference to facilitate problem-solving. We empirically find that more complicated goals can be derived from these fundamental manipulations.
We then define the standard CoM data structure to streamline the subsequent data construction and validation process. Given a question $Q$ about an initial input image $I_{0}$, a VLM equipped with chain of manipulations mechanism solves the problem to achieve final answer as $\textit{VLM}_{\boldsymbol{\varsigma}}(A,C|I_{0},Q)$, where $\boldsymbol{\varsigma}$ refers to the reasoning chain with evidence,
$\begin{split}\boldsymbol{\varsigma}&=(step_{1},step_{2},...)\\ step_{i}&=(f_{i},c_{i}),\quad\ f_{i}\in\mathcal{M}\end{split}$ (1)
where $C=(c_{i},c_{2},...,c_{|C|})$ refers to the free-form textual descriptions incorporating manipulation names $f_{i}$ and corresponding results from utilizing $f_{i}$. This definition explicitly declares the symbolic execution process, while also being compatible with linguistic reasoning steps. Based on this definition, we can clearly construct standard CoM samples that incorporating the manipulation executions and linguistic steps with evidence. After the data construction, we can utilize a simple method to convert the standard CoM samples to the compatible VQA samples.
@@ 3 Data Collection
Figure 4: A cascading data generation pipeline that automatically produces standard CoM samples. Given an original VQA sample, the linguistic annotator (LLMs) taught with usage of manipulations (prompt) is first asked to provide solving steps for the question $\mathcal{Q}$, and the visual foundational models (VFMs) are then engaged to replace the manipulations results, followed by a final traversal on the tree branched by the possible manipulation results to find positive paths terminating to the answer $\mathcal{A}$.
In this section, we first introduces the automated data generation pipeline (illustrated in Figure 4), that employs reliable LLMs as linguistic annotators and VFMs as the visual annotators to produce error-free CoM samples upon prevalent VQA corpus, and then present the manual annotation of high-quality CoM samples for the challenging graphical mathematical problems.
@@ 3.1 Automated Data Generation
Given a general corpus $\mathcal{D}=\{(I,Q,A)\}$ consisting of triplet samples of images with corresponding visual question-answer pairs, our automated data generation pipeline consists of a linguistic annotator and several visual annotators according to the manipulations. For a question $Q$ in each sample, we first engage the linguistic annotator to generate manipulations-assisted solving steps with the CoM format $(f_{i},c_{i})$, where the corresponding results of the instantiated manipulation executions are set with variables as placeholders. In this paper, we adopt GPT-4 (OpenAI, 2023a), a large language model with reliable language understanding and generation abilities as the linguistic annotator. We design a comprehensive prompt including the task requirements, usage of manipulations, and output data format, and further manually annotate 5 demonstrations for a stable generation. The detailed implementations are available at Appendix C.4.
We then employ essential visual annotators to supply the results of manipulations requested in the solving steps by exactly performing the corresponding manipulations. By empirically analyzing the manipulations from both predefined set and newly created ones (refers to Appendix C.3 for a detailed statistics), we reveal the Grounding and OCR are two fundamental manipulations, and most of the others can be consequently derived (e.g., CropZoomIn along a region of box, Counting upon recognized boxes, and Calculate for the recognized formula). Therefore, we employ two visual foundational models, GroundingDINO (Liu et al., 2023c) and PaddleOCR (Du et al., 2020), and develop the implementations of these manipulations [^1] We simply implement the CropZoomIn referring to human behaviors with a local code interpreter.. The execution of the manipulations will transform the sequential reasoning steps into a tree $\mathcal{T}$, as the input of current manipulation $f_{1}(x_{a})$ may rely on one of the multiple results of previous manipulation $f_{2}\rightarrow(x_{b},x_{c})$, i.e., $x_{a}$ rely on $x_{b}$ (e.g., step 2 for finding pillars in Figure 5). We then perform a traversal on each produced tree with Depth First Search (DFS) to find all positive paths $\{\mathcal{P}_{i}|\mathcal{P}_{i}\in\mathcal{T},i=1,2,...\}$ that can terminate with the final answer $A$ from the result of the last manipulation. Based on this method, the generated CoM samples with positive paths are guaranteed to be error-free. We implement this pipeline on 3 existing datasets that require detailed recognition or objects counting, TextVQA (Singh et al., 2019), ST-VQA (Biten et al., 2019), and TDIUC (Shrestha et al., 2019), to build 70K CoM samples  [^2] The success rate of GPT-4 to achieve the positive paths is 0.3555.. The designed prompt, a generated example with linguistic and visual results, and detailed algorithm illustration are available at AppendixC.1.
@@ 3.2 Human Annotation
The analysis from Fig.1 of AlphaGeometry (Trinh et al., 2024) shows that outputting auxiliary lines in linguistic reasoning process helps LLMs to solve complex geometry problems. Benefiting from the expressive capability of CoM structure, we have also manually annotated high-quality CoM samples for the graphical mathematical problems to facilitate VLMs in solving this challenging scenario. Similar to the automated pipeline, we engage 10 human experts as the linguistic annotators and visual annotators, where each expert is asked to annotate the linguistic solving steps and the use of manipulations, as well as the results of manipulations on images. We perform this annotation on the MathVista (Lu et al., 2023) and ChartQA (Masry et al., 2022), which include geometric and chart math problems, resulting in the collection of 6K high-quality CoM math samples.
Finally, we adapt the CoM samples to be compatible with VQA-style training samples. For each CoM sample including $n$ images from manipulations outputs $(I_{0},Q,C_{0},I_{1},C_{1},...,I_{n},A)$, we convert it into a multi-turn VQA sample segmented by the images $[(I_{0},Q,C_{0}),(I_{1},\bar{Q},C_{1}),...,(I_{n},\bar{Q},A)]$, where $C_{i}$ represents the intermediate steps between $I_{i}$ and $I_{i+1}$, and $\bar{Q}$ is a simple prompt asking model to answer question based on history. This transformation converts CoM samples into multi-turn VQA samples that are compatible with existing VLMs training data. The detailed statistics of the data generation are available at Appendix C.3.
@@ 4.1 Architecture
We use the same model architecture as CogVLM (Wang et al., 2023b), a general VLM approach that involves four fundamental components: (1) a Visual Encoder, (2) an MLP Adapter, (3) an LLM Backbone, and (4) a Visual Expert Module, for a reliable multimodal understanding. Concretely, the pre-trained EVA2-CLIP-E (Sun et al., 2023a) with 4B parameters and Vicuna-7B-v1.5 (Chiang et al., 2023) are adopted as the visual encoder and LLM backbone, respectively. A two-layer MLP (SwiGLU (Shazeer, 2020)) is further engaged to map the output of the visual encoder into the linguistic space of the LLM backbone. The visual expert module adds the vision-specific weights into the attention layer and feed-forward layer of each block in the LLM backbone, resulting in a total of 6.5B additional parameters for the deep fusion of modalities.
Based on this general architecture, we develop a memory-based multi-turn multi-image VLM approach. Specifically, for a multi-turn VQA sample $[(I_{t},Q_{t},A_{t})|t=1,2,...]$, where $A_{t}$ refers to $C_{t}$ in CoM, we keep the accumulated KV memories of each layer in the LLM backbone throughout these turns. And at each turn $t$ in training and inference, we calculate the attention function $att$ as:
$\begin{split}att(\boldsymbol{X})&=softmax(\frac{\boldsymbol{Q}_{t}\boldsymbol{K}_{t}^{\prime T}}{\sqrt{d}})\boldsymbol{V}^{\prime}_{t}\\ \boldsymbol{K}^{\prime}_{t}&=\text{trunc}(\text{concat}(\boldsymbol{K}_{0},\boldsymbol{K}_{1},...,\boldsymbol{K}_{t}))\\ \boldsymbol{V}^{\prime}_{t}&=\text{trunc}(\text{concat}(\boldsymbol{V}_{0},\boldsymbol{V}_{1},...,\boldsymbol{V}_{t}))\end{split}$ (2)
where $\boldsymbol{Q}_{t}\in\mathbb{R}^{s\times d}$ is query representation of current layer, and the $\boldsymbol{K}^{\prime}_{t},\boldsymbol{V}^{\prime}_{t}\in\mathbb{R}^{(s\times t)\times d}$ refer to the concatenation of accumulated representations and will be further truncated if the sequence length $s\times t$ is greater than a predefined threshold. At $t>0$, the new image $I_{t}$ will be cropped from $I_{t-1}$ and amplified with the Bicubic Interpolation (Keys, 1981).

### v3

# CogCoM: A Visual Language Model with Chain-of-Manipulations Reasoning

@@ Abstract
Vision-Language Models (VLMs) have shown broad effectiveness due to extensive training that aligns visual inputs with corresponding language responses. However, this conclusive alignment training causes models to overlook essential visual reasoning, leading to failures in handling detailed visual tasks and producing unfaithful responses. Drawing inspiration from human cognition in solving visual problems (e.g., marking, zoom in), this paper introduces Chain of Manipulations, a mechanism that enables VLMs to tackle problems step-by-step with evidence. After training, models can solve various visual problems by eliciting intrinsic manipulations (e.g., grounding, zoom in) with results (e.g., boxes, image) actively without relying on external tools, while also allowing users to trace error causes. In this paper, we study the comprehensive methodology that includes: (1) a flexible design of manipulations based on extensive analysis, (2) an efficient automated data generation pipeline, (3) a compatible VLM architecture capable of multi-turn, multi-image, and (4) a model training process for versatile capabilities. With the design, we also manually annotate 6K high-quality samples for challenging graphical mathematical problems. Our trained model, CogCoM, equipped with this mechanism and 17B parameters, achieves SOTA performance across 9 benchmarks in 4 categories, demonstrating its effectiveness while maintaining interpretability. Code, model, and data are available at https://github.com/THUDM/CogCoM.
Figure 1: CogCoM solves various visual problems with Chain of Manipulations (CoM) reasoning, which generates evidential and explainable steps without relying on external tools.
@@ 1 Introduction
Benefiting from the advantage of Large Language Models (LLMs) in broad world knowledge, large Vision-Language Models (VLMs) (Alayrac et al., 2022; Wang et al., 2023a) that are further trained to understand visual inputs have shown strong capabilities across a wide range of multimodal scenarios, such as visual question answering (Liu et al., 2023b), visual grounding (Peng et al., 2023), and optical character recognition (Zhang et al., 2023b). Research employing VLMs as foundation models (Bai et al., 2023; Sun et al., 2023b; Wang et al., 2023a) typically involves two main training stages, where the first stage develops intrinsic visual understanding through exposure to massive image-caption pairs, while the second stage endows the models with problem-solving abilities via the instruction tuning.
@@
However, existing tuning methods train models to provide conclusive responses to instructions based on visual inputs, causing them to overlook essential intermediate visual reasoning. This often leads to failures in handling visual tasks requiring subtle details, producing unfaithful responses, and even generating hallucinations. For example in the left subplot of Figure 2, we test the top-performing model, CogVLM (Wang et al., 2023a) on the details in the image (i.e., text written on a pillar), and it directly responds an incorrect answer (i.e., NO SMOKING), most likely due to bias to visual or linguistic priors (i.e., typical office scenes with a pillar). The lack of essential visual reasoning about the visual scene can result in a hasty response (Hwang et al., 2023).
@@
Humans solve problems regarding visual details by marking or processing the given images for convenience and precision, which we refer to as manipulations. For example, we identify targets by sequentially locating references, and focus on subtle details by zooming into the corresponding region. Most VLMs develop numerous intrinsic capabilities (e.g., grounding boxes, recognizing text) during the first stage of training. By further imitating the fundamental human behaviours (e.g., cropping, zooming in), models have the potential to perform this cognitive reasoning process. Three major obstacles in enabling VLMs to perform such reasoning are (1) the flexible definitions of manipulations that cover most visual problems, (2) an efficient data collection pipeline capable of producing abundant training data, and (3) a multi-turn multi-image structure that is compatible with existing models.
@@
Inspired by the human cognition in solving visual problems, we introduce Chain of Manipulations (CoM), a mechanism that enables VLMs to solve problems step-by-step with evidence, with each step potentially involving a manipulation on the visual input and its corresponding result, both generated by the model to facilitate the success and fidelity. This paper studies a comprehensive methodology, including manipulations design, data collection, model architecture, and training process for developing general VLMs with this mechanism. Based on pilot experiments, we first formally design 6 atomic manipulations, that are capable of addressing diverse visual problems. Next, we propose a cascading data generation pipeline that leverages reliable large language models (LLMs, serving as linguistic annotators) and visual foundational models (VFMs, serving as visual annotators), to automatically generate abundant training data. We collect 70K training samples involving visual reasoning chains using this pipeline. We then devise a multi-turn multi-image model architecture that is compatible with typical VLMs. Based on a data recipe incorporating the curated corpus, we finally train a general VLM equipped with CoM reasoning mechanism, named CogCoM, which possesses capabilities of Chat, Captioning, Grounding and Reasoning. Additionally, benefiting from the expressive capability of the proposed mechanism, we manually annotated 7K high-quality graphical mathematical samples, each accompanied by a CoM reasoning process, to further advance VLM research in solving challenging mathematical tasks.
@@
We conduct extensive experiments on 9 benchmarks from 4 categories, including TextVQA (Singh et al., 2019), ST-VQA (Biten et al., 2019), TallyVQA (Acharya et al., 2019), and GQA Hudson & Manning (2019) for detailed visual question answering, RefCOCO (Yu et al., 2016), RefCOCO+(Yu et al., 2016), and RefCOCOg (Mao et al., 2016) for visual grounding, POPE (Li et al., 2023d) for hallucination validation, and MM-Vet (Yu et al., 2023b) for general multimodal ability. Our model achieves up to 9.0 and 1.09 accuracy improvement on the detailed VQA and grounding benchmarks, respectively, and the superior performance on the general multimodal benchmark. The results demonstrate the effectiveness of the mechanism while preserving the interpretability of outputs.
@@ 2 Terminology
We first conduct pilot experiments to explore potential atomic manipulations that can address diverse visual problems.
@@
Specifically, given a question about an image, we prompt the advanced large language model, GPT-4, to generate solving steps by optionally utilizing possible actions on the image to facilitate problem-solving. We conduct this experiment on 170K questions from TextVQA, a dataset requiring detailed reasoning and recognition on images. To ensure stability, we manually write 4 demonstrations as priors. The detailed statistics are available in Appendix E.3.
@@
Based on the observation, we formally predefine a set of 6 atomic manipulations, which can either be developed during pre-training or learned through fine-tuning by imitating human behaviors: $\mathcal{M}\subseteq$$\textit{OCR}(tgt)\rightarrow txt$, $\textit{Grounding}(tgt)\rightarrow bbx$, $\textit{Counting}(tgt)\rightarrow num$, $\textit{Calculate}(tgt)\rightarrow num$, $\textit{CropZoomIn}(bbx,x)\rightarrow img$, $\textit{Line}(pts)\rightarrow img$}, where the corresponding parameters or results $tgt,txt,bbx,num,x,img,pts$ refer to the target description, texts, bounding boxes, numbers, zoom ratio, image, and points, respectively. In addition to the predefined manipulations, we also allow models to create new manipulations during inference to facilitate problem-solving. We empirically find that more complex goals can be derived from these fundamental manipulations.
We then define the standard CoM data structure to streamline the subsequent data construction and validation process. Given a question $Q$ about an initial input image $I_{0}$, a VLM equipped with chain of manipulations mechanism solves the problem to achieve final answer as $\textit{VLM}_{\bm{\varsigma}}(A,C|I_{0},Q)$, where $\bm{\varsigma}$ refers to the reasoning chain with evidence,
$\begin{split}\bm{\varsigma}&=(step_{1},step_{2},...)\\ step_{i}&=(f_{i},c_{i}),\quad\ f_{i}\in\mathcal{M}\end{split}$ (1)
where $C=(c_{i},c_{2},...,c_{|C|})$ refers to the free-form textual descriptions incorporating manipulation names $f_{i}$ and corresponding results by utilizing $f_{i}$. This definition explicitly declares the symbolic execution process, while remaining compatible with linguistic reasoning steps. Based on this definition, we can clearly construct standard CoM samples that incorporate the manipulation executions and linguistic steps with evidence. After the data construction, we can utilize a simple method to convert the standard CoM samples to the format of compatible VQA structure.
@@ 3 Data Collection
Figure 4: A cascading data generation pipeline that automatically produces standard CoM samples. Given an original VQA sample, the linguistic annotator (LLM) taught with usage of manipulations (prompt) is first asked to provide solving steps for the question $\mathcal{Q}$, and the visual foundational models (VFMs) are then engaged to replace the manipulations results, followed by a final traversal on the tree branched by the possible manipulation results to find positive paths terminating to the answer $\mathcal{A}$.
In this section, we first introduce the automated data generation pipeline (illustrated in Figure 4), that employs reliable LLMs as linguistic annotators and VFMs as the visual annotators to produce CoM samples upon prevalent VQA corpus, and then present the manual annotation of high-quality CoM samples for the challenging graphical mathematical problems.
@@ 3.1 Automated Data Generation
Given a general corpus $\mathcal{D}=\{(I,Q,A)\}$ consisting of triplet samples of images with corresponding visual question-answer pairs, our automated data generation pipeline consists of a linguistic annotator and several visual annotators according to the manipulations. For a question $Q$ in each sample, we first engage the linguistic annotator to generate manipulations-assisted solving steps with the CoM format $(f_{i},c_{i})$, where the corresponding results of the instantiated manipulation executions are set with variables as placeholders. In this paper, we adopt GPT-4 (OpenAI, 2023a), a large language model with reliable language understanding and generation abilities as the linguistic annotator. We design a comprehensive prompt including the task requirements, usage of manipulations, and output data format, and further manually annotate 5 demonstrations for a stable generation. The detailed implementations are available in Appendix E.4.
We then employ essential visual annotators to supply the results of manipulations requested in the solving steps by exactly performing the corresponding manipulations. By empirically analyzing the manipulations from both predefined set and newly created ones (refers to Appendix E.3 for a detailed statistics), we reveal the Grounding and OCR are two fundamental manipulations, and most of the others can be consequently derived (e.g., CropZoomIn along a region of box, Counting upon recognized boxes, and Calculate for the recognized formula). Therefore, we employ two visual fundamental models, GroundingDINO (Liu et al., 2023c) and PaddleOCR (Du et al., 2020), and develop the implementations of these manipulations [^1] We simply implement the CropZoomIn referring to human behaviors with a local code interpreter.. The execution of the manipulations will transform the sequential reasoning steps into a tree $\mathcal{T}$, as the input of current manipulation $f_{1}(x_{a})$ may rely on one of the multiple results of previous manipulations $f_{2}\rightarrow(x_{b},x_{c})$, i.e., $x_{a}$ rely on $x_{b}$ (e.g., step 2 for finding pillars in Figure 5). We then perform a Depth First Search (DFS) traversal on each produced tree to find all positive paths $\{\mathcal{P}_{i}|\mathcal{P}_{i}\in\mathcal{T},i=1,2,...\}$, that terminates with the golden answer $A$ as the result of the last manipulation. Based on this recursively searching method, most of the generated positive paths are guaranteed to be error-free. We implement this pipeline on 3 existing datasets that require detailed recognition or counting, TextVQA (Singh et al., 2019), ST-VQA (Biten et al., 2019), and TDIUC (Shrestha et al., 2019), to build 70K CoM samples  [^2] The success rate of GPT-4 to achieve the positive paths is 0.3555.. The prompt design, an example with linguistic and visual results, and algorithm are available in AppendixE.1.
@@ 3.2 Human Annotation
The analysis from Fig.1 of AlphaGeometry (Trinh et al., 2024) shows that outputting auxiliary lines in linguistic reasoning process helps LLMs to solve complex geometry problems. Benefiting from the expressive capability of CoM structure, we have also manually annotated high-quality CoM samples for the graphical mathematical problems to facilitate VLMs in solving this challenging scenario. Similar to the automated pipeline, we engage 10 human experts as the linguistic annotators and visual annotators, where each expert is asked to annotate the linguistic solving steps and the use of manipulations, as well as the results of manipulations on images. We perform this annotation on the MathVista (Lu et al., 2023) and ChartQA (Masry et al., 2022), which include geometric and chart math problems, resulting in the collection of 7K high-quality CoM math samples.
Finally, we adapt the CoM samples to be compatible with VQA-style training samples. For each CoM sample including $n$ images with manipulations outputs $(I_{0},Q,C_{0},I_{1},C_{1},...,I_{n},A)$, we convert it into a multi-turn VQA sample segmented by the images $[(I_{0},Q,C_{0}),(I_{1},\bar{Q},C_{1}),...,(I_{n},\bar{Q},A)]$, where $C_{i}$ represents the intermediate steps between $I_{i}$ and $I_{i+1}$, and $\bar{Q}$ is a simple prompt asking model to answer question based on history. This transformation converts CoM samples into multi-turn VQA samples that are compatible with existing VLMs training corpus. The detailed statistics of the data generation are available at Appendix E.3.
@@ 4.1 Architecture
We use the same model architecture as CogVLM (Wang et al., 2023a), a general VLM approach that involves four fundamental components: (1) a Visual Encoder, (2) an MLP Adapter, (3) an LLM Backbone, and (4) a Visual Expert Module, for a reliable multimodal understanding. Concretely, the pre-trained EVA2-CLIP-E (Sun et al., 2023a) with 4B parameters and Vicuna-7B-v1.5 (Chiang et al., 2023) are adopted as the visual encoder and LLM backbone, respectively. A two-layer MLP (SwiGLU (Shazeer, 2020)) is further engaged to map the output of the visual encoder into the linguistic space of the LLM backbone. The visual expert module adds the vision-specific weights into the attention layer and feed-forward layer of each block in the LLM backbone, resulting in a total of 6.5B additional parameters for the deep fusion of modalities.
Based on this general architecture, we develop a memory-based multi-turn multi-image VLM architecture. Specifically, for a multi-turn VQA sample $[(I_{t},Q_{t},A_{t})|t=1,2,...]$, where $A_{t}$ refers to $C_{t}$ in CoM, we keep the accumulated KV memories of each layer in the LLM backbone throughout these turns. And at each turn $t$ in training and inference, we calculate the attention function $att$ as:
$\begin{split}att(\bm{X})&=softmax(\frac{\bm{Q}_{t}\bm{K}_{t}^{\prime T}}{\sqrt{d}})\bm{V}^{\prime}_{t}\\ \bm{K}^{\prime}_{t}&=\text{trunc}(\text{concat}(\bm{K}_{0},\bm{K}_{1},...,\bm{K}_{t}))\\ \bm{V}^{\prime}_{t}&=\text{trunc}(\text{concat}(\bm{V}_{0},\bm{V}_{1},...,\bm{V}_{t}))\end{split}$ (2)
where $\bm{Q}_{t}\in\mathbb{R}^{s\times d}$ is query representation of current layer, and the $\bm{K}^{\prime}_{t},\bm{V}^{\prime}_{t}\in\mathbb{R}^{(s\times t)\times d}$ refer to the concatenation of accumulated representations and will be further truncated if the sequence length $s\times t$ is greater than a predefined threshold. At $t>0$, the new image $I_{t}$ will be cropped from $I_{t-1}$ and amplified with the Bicubic Interpolation (Keys, 1981).